# Real Noncommutative Convexity II: Extremality and nc convex functions

### v2

@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
(See [33] for an alternative take on some of the facts concerning $\mathbb{C}_{r}$ below.) Note that $\mathbb{H},\mathbb{C}_{r}$ each have only one real *-irrep up to unitary equivalence. This follows from [29, Proposition 5.3.7 (1)] because these algebras have only one real state, hence only one pure real state. Thus e.g. all irreps of $\mathbb{H}$ are 4 dimensional and unitarily equivalent to the standard matrix form of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. Thus the real nc Choquet boundary of $\mathbb{C}_{r}$ is the set of elements $U^{T}\,c\,U$ where $c:\mathbb{C}\to M_{2}(\mathbb{R})$ is as usual, and $U$ is a rotation or reflection matrix. On the other hand, the complexification of $\mathbb{C}$ has two irreducible representations, and the complex nc Choquet boundary consists of two points. This shows that in general for a general nc convex set $K$, there need be no nice relation between the real Choquet boundary of $A(K)$ and the complex Choquet boundary in $K_{c}$. However the nc Shilov boundaries (which can be identified with the $C^{*}$-envelopes) are of course closely related. The noncommutative state space of $\mathbb{C}_{r}$ is identifiable with the nc convex set ${\mathcal{C}}{\mathcal{S}}$ of skew symmetric real matrices of norm $\leq 1$. One can show by functional calculus that these are exactly the matrices $\varphi(i)$ for real ucp $\varphi:\mathbb{C}\to M_{n}(\mathbb{R})$. The noncommutative state space of $\mathbb{H}$ was discussed earlier and in [8]. These two sets coincide with the nc Shilov boundary, and are the nc closed convex hull of the nc Choquet boundary above, by Theorem 2.21. For the $\mathbb{C}$ example this reveals an interesting relation between the skew symmetric real matrices of norm $\leq 1$, and the single matrix $c(i)=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$. Namely, they are the noncommutative convex hull of $c(i)$: that is, ${\mathcal{C}}{\mathcal{S}}={\rm nconv}(c(i))$. All of the $C^{*}$-extreme points of this set are computed in [33], and these turn out to correspond to the *-representations. So they are exactly the maximal elements by Proposition 2.12. The nc extreme points will be those *-representations which are irreducible, and we said that these are unitarily equivalent to the one in Example 3.5. The complexification of the nc convex set ${\mathcal{C}}{\mathcal{S}}$ turns out to be the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. See Example 3.6 for more details. It would be quite interesting to perform a similar analysis to the above with the quaternions.
@@
In this case $C^{*}_{\rm min}(V)$ is a finite dimensional real $C^{*}$-algebra. Hence by [29, Theorem 5.7.1] we may assume that $C^{*}_{\rm min}(V)=\oplus_{k=1}^{N}\,M_{n_{k}}(D_{k})$ where $D_{k}$ is $\mathbb{R},\mathbb{C}$ or $\mathbb{H}$. The real irreps of this sum are up to unitary equivalence the obvious ones $\pi_{k}$ associated with the summands. For example, corresponding to the canonical irreps $M_{n}(\mathbb{C}_{r})\to B(\mathbb{R}^{2n})$ and $M_{n}(\mathbb{H})\to B(\mathbb{R}^{4n})$. This is easy to see using the fact that $\mathbb{H},\mathbb{C}_{r}$ each have only one real *-irrep up to unitary equivalence. We finish by an argument of Arveson: For any $j$, we have $\cap_{k\neq j}\,{\rm Ker}(\pi_{k})\neq(0)$, obviously. However if $\pi_{j}$ is not a boundary representation then $\cap_{k\neq j}\,{\rm Ker}(\pi_{k})=(0)$ since this collection includes (up to unitary equivalence) all boundary representations, and there are sufficiently many boundary representations. ∎
@@
The intersection in $C^{*}_{\rm min}(V)$ of the kernels of the boundary representations is $(0)$ (e.g. see the discussion above Theorem 2.15). Hence the corollary follows by the real case of basic facts about the topology of the spectrum of a $C^{*}$-algebra. Some of these basic facts may be found in [29] (e.g. Section 5.3 there), and in places in the proof of Theorem 5.3 in [9]), or have already been discussed in the proof of Theorem 2.20 above. Any other aspects of the real case of classical results about the spectrum $\hat{A}$ in early chapters of [17] which we need here are easy to check and left to the interested reader. ∎
@@ 3. Complexification of extremals
In this section we examine the connections between extreme, pure, and maximal points in the real and complex case, that is, how the these notions interact with the complexification of a nc convex set.
Classically pure (or equivalently, classical extreme [29, Proposition 5.3.5 (1)]) elements in $K_{1}$, are clearly just the nc pure states (in both the real and complex case). For example, if a state $\varphi$ is nc pure and is a convex combination of other states, then it is easy to see that these states must be $\varphi$. However they are not necessarily nc extreme/maximal. An example is the obvious real pure state on $\mathbb{C}$, which is not maximal.
@@
For $K$ a real compact nc convex set, if $x\in K$ is pure (resp. extreme) in $K_{c}$ then $x$ is pure (resp. extreme) in $K$. The converse is false: pure (resp. extreme) in $K$ need not imply pure (resp. extreme) in $K_{c}$. In particular, in general
@@
If $x$ is extreme in $K_{c}$ then in addition each $x_{k}$ commutes with $\alpha_{k}\,\alpha_{k}^{\mathsf{T}}$.
@@
For $K$ a real compact nc convex set, if $x\in K_{n}$ is extreme in $K$ then $x$ is extreme in $K_{c}$ if and only if $x$ is complex-irreducible in $M_{n}(\mathbb{C})$, or equivalently every nonzero positive in $M_{n}(\mathbb{C})$ commuting with $x$ is a positive scalar multiple of $I_{n}$. In particular, $x\in K_{1}$ is nc extreme in $K$ if and only if $x+i0$ is nc extreme in $K_{c}$.
@@
Remark. The class of complex-irreducible extreme or pure real matrix states is in general not big enough to norm an operator system or even a real $C^{*}$-algebra. Indeed for $\mathbb{C}_{r}$ there are no complex-irreducible extreme real matrix states (as is said elsewhere the irreducible real matrix states of $\mathbb{C}_{r}$ are all unitarily equivalent to the one in Example 3.5, and none of these are complex-irreducible).
Lemma 3.3.
For $K$ a real compact nc convex set, if $x\in K_{c}$ then $x$ is complex maximal in $K_{c}$ if and only if it is real maximal in $K_{c}$.
Proof.
Certainly every complex maximal is real maximal. For the other, suppose $x\in(K_{c})_{n}$ is real maximal. Write $x=\alpha^{*}y\alpha$ for $y\in(K_{c})_{l}$ and $\alpha\in M_{l,n}(\mathbb{C})$ an isometry. Taking $c$ of both sides gives $c(x)=c(\alpha)^{\mathsf{T}}c(y)c(\alpha)$. Compressing both sides by $v=[1_{n}\;0]^{\mathsf{T}}$ gives
$x=(c(\alpha)v)^{\mathsf{T}}c(y)(c(\alpha)v).$
Now $c(\alpha)v$ is a real isometry and so by the fact that $x$ is real maximal we get that $c(y)$ commutes with the projection
$P=c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}.$
Therefore, $Pc(y)=c(y)P$ and adjoining $u_{n}$ to this gives
$u_{l}^{*}c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}c(y)u_{l}=u_{l}^{*}c(y)c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}u_{l}.$
As in [8, Lemma 3.8] for $c(v)$ in either $M_{2m,2k}(\mathbb{C})$ or $M_{2m,2k}(K)$ we have that $u_{m}^{*}c(v)=vu_{k}^{*}$ and similarly $c(v)u_{k}=u_{m}v$. Therefore, the left hand side of the above equality turns into
$u_{l}^{*}c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}c(y)u_{l}$ $=\frac{1}{2}\alpha\begin{bmatrix}1_{n}&i1_{n}\end{bmatrix}\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}\begin{bmatrix}1_{n}\\ -i1_{n}\end{bmatrix}\alpha^{*}y$
$=\frac{1}{2}\alpha\alpha^{*}y,$
$\alpha\alpha^{*}y=y\alpha\alpha^{*}.$
So, $y$ decomposes with respect to the range of $\alpha$ as a direct sum. ∎
Remark. For an element $x$ of a general complex compact nc convex set $K$, we do not know if $x$ is complex maximal in $K$ if it is real maximal in $K$. However we just proved that this does hold if $K$ is a complexification, and the next proposition is a partial result in this direction. It is worth remarking that any complex compact nc convex set that is not a complexification, is probably in a sense not easy to describe. Indeed such would give a complex operator system that is not an operator system complexification, and there are no known simple examples of this at this date (we thank Chris Phillips for a recent conversation on this point). For $C^{*}$-algebras this is equivalent to the hard problem of the existence of a real form (see e.g. first page or two of [37]). There are not even any known simple low dimensional examples of complex Banach spaces that are not (isometric) complexifications.
@@
Suppose that $\varphi=\alpha^{*}\psi\alpha$ for $\psi\in{\rm UCP}_{\mathbb{C}}(B,M_{m}(\mathbb{C}))$ and real isometry $\alpha\in M_{m,n}(\mathbb{R})\subset M_{m,n}(\mathbb{C})$. View $\psi$ as in ${\rm UCP}_{\mathbb{R}}(B,M_{2m}(\mathbb{R}))$ and $c_{n}\circ\varphi$ in ${\rm UCP}_{\mathbb{R}}(B,M_{2n}(\mathbb{R}))$. Then $c_{n}\circ\varphi=c(\alpha+i0)^{\mathsf{T}}\,c_{m}\circ\psi\,c(\alpha+i0)$. Since $c_{n}\circ\pi$ is real maximal in $\mathrm{ncS}_{\mathbb{R}}(B)$ we have that $c(\alpha+i0)c(\alpha+i0)^{\mathsf{T}}=c(\alpha\alpha^{*}+i0)$ commutes with $c_{m}\circ\psi$. However this forces $\alpha\alpha^{*}$ to commute with $\psi$ as desired. ∎
If $x$ is pure in $K$ then it need not be real nor complex pure in $K_{c}$ (see Example 3.6). However if $x$ is extreme in $K$ then it is real extreme (hence real pure) in $K_{c}$, by a small variant of the proof of Corollary 3.2.
@@
We exhibit a simple example of an extreme $x\in K_{2}$ which is not extreme in $K_{c}$. Indeed the criterion in Corollary 3.2 is not true in general for $n=2$. For $A=\mathbb{C}$ consider the usual representation $c:A\to M_{2}(\mathbb{R})$. This is a real irreducible *-representation, hence is extreme and pure by Proposition 2.12 (2). However if we take
@@
Consider the noncommutative convex set
[…]
From Theorem $3.1$ in [8], we know that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\mathrm{ncconv}_{\mathbb{R}}\{c(i)\}\}=\mathrm{ncconv}_{\mathbb{C}}\{c(i)\}$. A compression by $u$ shows $\pm i\in K_{c}$ so that $\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}\subseteq K_{c}$. Also,
@@
so that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}$. Let $L=\mathrm{ncconv}_{\mathbb{C}}\{\pm 1\}$, the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. Because $\pm i\in i\cdot L$ we have that $K_{c}\subseteq iL$. Conversely, $\pm 1\in-i\cdot\mathrm{ncconv}\{\pm i\}$ so that $L\subseteq-iK_{c}$. This implies $iL\subseteq K_{c}$ and thus $iL=K_{c}$. However, $L=\mathrm{ncconv}\{\pm 1\}=\bigsqcup_{n}[-1,1]_{n}$ by Example $4.1$ in [40]. Also, $iL\cong L$ because multiplication by $i$ is a nc affine homeomorphism. Now, $L$ is compact, implying that $K_{c}$ is too, and thus $K$ is compact by Theorem 3.1. We have that $\pm 1\in L$ are the only nc extreme points also by Example $4.1$ in [40], so that $\pm c(i)$ are the only possible extreme points in $K$, and they are indeed extreme. In Example 2.22 we observed that $K=\mathrm{ncS}(\mathbb{C}_{r})$, and equals the skew matrix set ${\mathcal{C}}{\mathcal{S}}$. So $A(K)\cong\mathbb{C}_{\mathbb{R}}$ as operator systems.
@@
Let $I=\mathrm{ncconv}_{\mathbb{R}}\{\pm 1\}$. This equals the compact real operator interval $\bigsqcup_{n}[-1,1]_{n}$. To see this, $\subseteq$ is obvious, and for the converse we observe that $\mathrm{ncconv}_{\mathbb{R}}\{\pm 1\}=\mathrm{ncconv}_{\mathbb{R}}\{[-1,1]\}$. Elements of the compact operator interval can all be written as the diagonalization $U^{\mathsf{T}}DU$ for $D$ the direct sum of elements in the interval [1,1][-1,1]. Thus, all elements in the operator interval are in $I$. So $I$ is the compact real operator interval. Elements of $A(I)$ are determined by where they send $\pm 1$. Therefore, $A(I)$ is a two dimensional real operator system, indeed is a real $C^{*}$-algebra *-isomorphic to $\ell^{\infty}_{2}(\mathbb{R})$. The isomorphism is the map $\theta:A(I)\to\ell^{\infty}_{2}(\mathbb{R})$ taking $f\in A(I)$ to (f(1),f(1))(f(-1),f(1)). Note that $A(I)$ has trivial involution, so all we need to check is that $\theta$ is multiplicative. However,
Because $\theta$ is bijective, this is a *-isomorphism. Although both $I$ and ${\mathcal{C}}{\mathcal{S}}=K$ have the same complexification they are not nc affine homeomorphic, indeed their associated operator systems are $\ell^{\infty}_{2}(\mathbb{R})$ and $\mathbb{C}_{r}$, which are quite different. Indeed although $I={\rm Min}([-1,1])$ in the language of [8, Section 6], $K$ is not a Min nc convex set. Indeed $A(K)=\mathbb{C}_{r}$ is not a real OMIN operator system: it has nontrivial involution (has a skew element) and is not a real function system in the technical sense of [9, 8]. Indeed we showed in [9, Section 10] that an operator system cannot be an OMIN if it has a skew element, and as we mentioned earlier a real $C^{*}$-algebra with no skew element is a real $C(X)$.
Note that $A(K)$ is not a $C^{*}$-algebra under the usual product and involution. If it were, then the map $\Xi:A(K)\to\mathbb{R}$ given by $\Xi(f)=f(0)$ is easily seen to be a *-homomorphism. Therefore, $0\in\mathrm{ncS}(A(K))$ would be a representation, and thus maximal. This is a contradiction.
@@
For $K$ a real compact nc convex set, if $x\in K_{c}$ has $c(x)$ pure (resp. extreme, maximal) in $K$ then $x$ is pure (resp. extreme, maximal) in $K_{c}$.

### v3

@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
(See [36] for an alternative take on, and proofs of, some of the facts concerning $\mathbb{C}_{r}$ below.) Note that $\mathbb{H},\mathbb{C}_{r}$ each have only one real *-irrep up to unitary equivalence. This follows from [32, Proposition 5.3.7 (1)] because these algebras have only one real state, hence only one pure real state. Thus e.g. all irreps of $\mathbb{H}$ are 4 dimensional and unitarily equivalent to the standard matrix form of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. Thus the real nc Choquet boundary of $\mathbb{C}_{r}$ is the set of elements $U^{T}\,c\,U$ where $c:\mathbb{C}\to M_{2}(\mathbb{R})$ is as usual, and $U$ is a rotation or reflection matrix. On the other hand, the complexification of $\mathbb{C}$ has two irreducible representations, and the complex nc Choquet boundary consists of two points. This shows that in general for a general nc convex set $K$, there need be no nice relation between the real Choquet boundary of $A(K)$ and the complex Choquet boundary in $K_{c}$. However the nc Shilov boundaries (which can be identified with the $C^{*}$-envelopes) are of course closely related.
The noncommutative state space of $\mathbb{C}_{r}$ is identifiable with the nc convex set ${\mathcal{C}}{\mathcal{S}}$ of skew symmetric real matrices of norm $\leq 1$. One can show by functional calculus that these are exactly the matrices $\varphi(i)$ for real ucp $\varphi:\mathbb{C}\to M_{n}(\mathbb{R})$. Indeed clearly such $\varphi(i)$ is in ${\mathcal{C}}{\mathcal{S}}$. Conversely if $A\in{\mathcal{C}}{\mathcal{S}}$ then $iA\in M_{n}(\mathbb{C})_{\rm sa}$, so that $iA=U^{*}{\rm diag}(d_{1},\cdots,d_{n})U$, where $d_{k}\in[-1,1]$ and $U$ is unitary. The selfadjoint map taking $s+it\in\mathbb{C}$ to $sI_{n}+tA=U^{*}{\rm diag}(s+itd_{1},\cdots,s+itd_{n})U$ in $M_{n}(\mathbb{R})$, is unital. Here $s,t\in\mathbb{R}$. It is also (completely) positive, since the unital scalar functionals $s+it\mapsto s+itd_{k}$ are ucp. The latter may be seen in several ways. For example it is equivalent to $c(s,t)\mapsto c(s,td)$ being ucp for $d\in[-1,1]$, which is clear for example since multiplication by $d$ on $\mathbb{R}$ induces a ucp map on the real Paulsen system of $\mathbb{R}$.
The nc state space of $\mathbb{H}$ was discussed earlier and in [8]. These two nc state spaces coincide with the nc Shilov boundary, and are the nc closed convex hull of the nc Choquet boundary above, by Theorem 2.21. For the $\mathbb{C}$ example this reveals an interesting relation between the skew symmetric real matrices of norm $\leq 1$, and the single matrix $c(i)=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$. Namely, they are the noncommutative convex hull of $c(i)$: that is, ${\mathcal{C}}{\mathcal{S}}={\rm nconv}(c(i))$.
All of the $C^{*}$-extreme points of ${\mathcal{C}}{\mathcal{S}}$ are computed in [36], and these turn out to correspond to the *-representations. So they are exactly the maximal elements by Proposition 2.12. The nc extreme points will be those *-representations which are irreducible, and we said at the start of this Example that these are unitarily equivalent to the usual representation $c:\mathbb{C}_{r}\to M_{2}(\mathbb{R})$ (see also Example 3.4). The complexification of the nc convex set ${\mathcal{C}}{\mathcal{S}}$ turns out to be the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. See Example 3.6 for more details. It would be quite interesting to perform a similar analysis to the above with the quaternions.
@@
In this case $C^{*}_{\rm min}(V)$ is a finite dimensional real $C^{*}$-algebra. Hence by [32, Theorem 5.7.1] we may assume that $C^{*}_{\rm min}(V)=\oplus_{k=1}^{N}\,M_{n_{k}}(D_{k})$ where $D_{k}$ is $\mathbb{R},\mathbb{C}$ or $\mathbb{H}$. The real irreps of this sum are up to unitary equivalence the obvious ones $\pi_{k}$ associated with the summands. For example, corresponding to the canonical irreps $M_{n}(\mathbb{C}_{r})\to B(\mathbb{R}^{2n})$ and $M_{n}(\mathbb{H})\to B(\mathbb{R}^{4n})$. This is easy to see using the fact that $\mathbb{H},\mathbb{C}_{r}$ each have only one real *-irrep up to unitary equivalence. We finish by an argument of Arveson: For any $j$, we have $\cap_{k\neq j}\,{\rm Ker}(\pi_{k})\neq(0)$, obviously. However if $\pi_{j}$ is not a boundary representation then $\cap_{k\neq j}\,{\rm Ker}(\pi_{k})=(0)$ since this collection includes (up to unitary equivalence) all boundary representations, and there are sufficiently many boundary representations. ∎
@@
The intersection in $C^{*}_{\rm min}(V)$ of the kernels of the boundary representations is $(0)$ (e.g. see the discussion above Theorem 2.15). Hence the corollary follows by the real case of basic facts about the topology of the spectrum of a $C^{*}$-algebra, such as [19, Theorem 3.4.10]. Some of these basic facts may be found in [32] (e.g. Section 5.3 there), and in places in the proof of Theorem 5.3 in [9]), or have already been discussed in the proof of Theorem 2.20 above. Any other aspects of the real case of classical results about the spectrum $\hat{A}$ in early chapters of [19] which we need here are easy to check and left to the interested reader. ∎
We end this section with a result continuing the earlier theme that ‘maximal’ elements behave perfectly with respect to “real versus complex distinctions” or complexification.
Theorem 2.25.
For $K$ a complex compact nc convex set and $x\in K$, then $x$ is complex maximal in $K$ if and only if it is real maximal in $K$.
Proof.
Certainly every complex maximal is real maximal. For the other direction, we may suppose that $0\in K_{1}$ by translation. Suppose that $x\in K_{n}$ is real maximal, and let $v=[I_{n}\;0]^{\mathsf{T}}$. Write $x=\alpha^{*}y\alpha$ for $y\in K_{m}$ and $\alpha\in M_{m,n}(\mathbb{C})$ an isometry. We may write $\alpha=\beta+i\lambda$ with $\beta,\lambda\in M_{m,n}(\mathbb{R})$. Taking $c$ of both sides of this and compressing both sides by $u_{n}$ gives $x=u_{n}^{*}c(\alpha)^{\mathsf{T}}c(y)c(\alpha)u_{n}$. Let $w=c((1-i)/\sqrt{2})$, a projection having $u$ as eigenvector with eigenvalue $\frac{1}{\sqrt{2}}$. It is easy to see that $c(\alpha)u_{n}=w_{m}c(\alpha)v$. Let $\gamma=c(\alpha)v=[\beta\;\lambda]^{\mathsf{T}}$, so that $c(\alpha)u_{n}=w_{m}\gamma$. We have $\gamma^{\mathsf{T}}\gamma=\beta^{\mathsf{T}}\beta+\lambda^{\mathsf{T}}\lambda=I_{n}$, since
$I_{n}=\alpha^{*}\alpha={\rm Re}\,(\beta^{\mathsf{T}}-i\lambda^{\mathsf{T}})(\beta+i\lambda)\,=\,\beta^{\mathsf{T}}\beta+\lambda^{\mathsf{T}}\lambda.$
Let $z=w_{m}^{*}c(y)w_{m}=w_{m}^{*}c(y)w_{m}+q0q\in K$, where $q$ is the complementary projection to $w_{n}$. Then
$x=u_{n}^{*}c(\alpha)^{\mathsf{T}}c(y)c(\alpha)u_{n}=\gamma^{\mathsf{T}}w_{m}^{*}c(y)w_{m}\gamma=\gamma^{\mathsf{T}}z\gamma.$
Since $x$ is real maximal we get that the projection
$P=\gamma\gamma^{\mathsf{T}}=c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}$
commutes with $z$. Therefore, $P(w_{m}c(y)w_{m})=(w_{m}c(y)w_{m})P$ and adjoining $u_{m}$ to this gives
$u_{m}^{*}c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}w_{m}c(y)w_{m}u_{m}=u_{m}^{*}w_{m}c(y)w_{m}c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}u_{m}.$
As in [8, Lemma 3.8] for $c(a)$ in either $M_{2m,2k}(\mathbb{C})$ or $M_{2m,2k}(K)$ we have that $u_{m}^{*}c(a)=au_{k}^{*}$ and similarly $c(a)u_{k}=u_{m}a$. Thus
$u_{m}^{*}\,c(\alpha)=\alpha\,u_{n}^{*}\;,\;\;\;\;\;\;c(\alpha)^{\mathsf{T}}\,u=u\,\alpha^{*}.$
Recall that $wu=\frac{1}{\sqrt{2}}u$, and that $c(y)=y\oplus y$ ‘commutes with’ block matrices whose blocks are scalar multiples of $I_{m}$, so that
$w_{m}c(y)w_{m}u_{m}=\frac{1}{\sqrt{2}}\,w_{m}c(y)u_{m}=\frac{1}{2}\,u_{m}\,y.$
Using these two displayed equations, the left hand side of the displayed equation above those, turns into
$u_{m}^{*}c(\alpha)\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}w_{m}c(y)w_{m}u_{m}$ $=\frac{1}{2}\,\alpha u_{n}^{*}\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}c(\alpha)^{\mathsf{T}}u_{m}y$
$=\frac{1}{2}\,\alpha u_{n}^{*}\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}u_{n}\alpha^{*}y$
$=\frac{1}{4}\,\alpha\begin{bmatrix}1_{n}&i1_{n}\end{bmatrix}\begin{bmatrix}1_{n}&0\\ 0&0\end{bmatrix}\begin{bmatrix}1_{n}\\ -i1_{n}\end{bmatrix}\alpha^{*}y$
$=\frac{1}{2}\,\alpha\alpha^{*}y,$
$\alpha\alpha^{*}y=y\alpha\alpha^{*}.$
So, $y$ decomposes with respect to the range of $\alpha$ as a direct sum. ∎
@@ 3. Complexification of extremals
In this section we examine the connections between nc extreme, pure, and maximal points in the real and complex case, that is, how these notions interact with the complexification of a nc convex set.
Classically pure (or equivalently, classical extreme [32, Proposition 5.3.5 (1)]) elements in $K_{1}$, are clearly just the nc pure states (in both the real and complex case). For example, if a state $\varphi$ is nc pure and is a convex combination of other states, then it is easy to see that these states must be $\varphi$. However they are not necessarily nc extreme/maximal. An example is the obvious real pure state on $\mathbb{C}$, which is not maximal.
@@
For $K$ a real compact nc convex set, if $x\in K$ is pure (resp. nc extreme) in $K_{c}$ then $x$ is pure (resp. nc extreme) in $K$. The converse is false: pure (resp. nc extreme) in $K$ need not imply pure (resp. nc extreme) in $K_{c}$. In particular, in general
@@
If $x$ is nc extreme in $K_{c}$ then in addition each $x_{k}$ commutes with $\alpha_{k}\,\alpha_{k}^{\mathsf{T}}$.
@@
For $K$ a real compact nc convex set, if $x\in K_{n}$ is nc extreme in $K$ then $x$ is nc extreme in $K_{c}$ if and only if $x$ is complex-irreducible in $M_{n}(\mathbb{C})$, or equivalently every nonzero positive in $M_{n}(\mathbb{C})$ commuting with $x$ is a positive scalar multiple of $I_{n}$. In particular, $x\in K_{1}$ is nc extreme in $K$ if and only if $x+i0$ is nc extreme in $K_{c}$.
@@
Remarks. 1) The class of complex-irreducible nc extreme or pure real matrix states is in general not big enough to norm an operator system or even a real $C^{*}$-algebra. Indeed for $\mathbb{C}_{r}$ there are no complex-irreducible extreme real matrix states (as is said elsewhere the irreducible real matrix states of $\mathbb{C}_{r}$ are all unitarily equivalent to the one in Example 3.4, and none of these are complex-irreducible).
2) It is worth remarking that any complex compact nc convex set that is not a complexification, is probably in a sense not easy to describe. Indeed such would give a complex operator system that is not an operator system complexification, and there are no known simple examples of this at this date (we thank Chris Phillips for a recent conversation on this point). For $C^{*}$-algebras this is equivalent to the hard problem of the existence of a real form (see e.g. first page or two of [40]). There are not even any known simple low dimensional examples of complex Banach spaces that are not (isometric) complexifications.
@@
Suppose that $\varphi=\alpha^{*}\psi\alpha$ for $\psi\in{\rm UCP}_{\mathbb{C}}(V,M_{m}(\mathbb{C}))$ and an isometry $\alpha\in M_{m,n}(\mathbb{C})$. View $\psi$ as in ${\rm UCP}_{\mathbb{R}}(V,M_{2m}(\mathbb{R}))$ and $c_{n}\circ\varphi$ in ${\rm UCP}_{\mathbb{R}}(V,M_{2n}(\mathbb{R}))$. Then $c_{n}\circ\varphi=c(\alpha)^{\mathsf{T}}\,(c_{m}\circ\psi)\,c(\alpha)$. Since $c_{n}\circ\varphi$ is real maximal in $\mathrm{ncS}_{\mathbb{R}}(V)$ we have that $c(\alpha)c(\alpha)^{\mathsf{T}}=c(\alpha\alpha^{*}+i0)$ commutes with $c_{m}\circ\psi$. However this forces $\alpha\alpha^{*}$ to commute with $\psi$ as desired. ∎
If $x$ is pure in $K$ then it need not be real nor complex pure in $K_{c}$ (for examples see 3.4 and 3.6, and Lemma 3.8). However if $x$ is nc extreme in $K$ then it is real nc extreme (hence real pure) in $K_{c}$, by a small variant of the proof of Corollary 3.2.
@@
We exhibit a simple example of a nc extreme $x\in K_{2}$ which is not nc extreme in $K_{c}$. Indeed the criterion in Corollary 3.2 is not true in general for $n=2$. For $A=\mathbb{C}$ consider the usual representation $c:A\to M_{2}(\mathbb{R})$. This is a real irreducible *-representation, hence is nc extreme and pure by Proposition 2.12 (2). However if we take
@@
For the next two examples, we use the real compact nc convex set
[…]
In Example 2.22 we observed that $K$ may be viewed as the nc state space of $\mathbb{C}_{r}$, and that $K$ is the skew matrix set $\mathcal{C}S$}. So $A(K)\cong\mathbb{C}_{r}$ as operator systems.
@@
Let $I=\mathrm{ncconv}_{\mathbb{R}}\{\pm 1\}$. This equals the compact real operator interval $\bigsqcup_{n}[-1,1]_{n}$. To see this, observe that $\subseteq$ is obvious and for the converse we observe that $\mathrm{ncconv}_{\mathbb{R}}\{\pm 1\}=\mathrm{ncconv}_{\mathbb{R}}\{[-1,1]\}$. Elements of the compact operator interval can all be written as the diagonalization $U^{\mathsf{T}}DU$ for $D$ the direct sum of elements in the interval [1,1][-1,1]. Thus, all elements in the operator interval are in $I$, which gives $\supseteq$.
Elements of $A(I)$ are determined by where they send $\pm 1$. Therefore, $A(I)$ is a two dimensional real operator system, indeed it is a real $C^{*}$-algebra *-isomorphic to $\ell^{\infty}_{2}(\mathbb{R})$. The isomorphism is the map $\theta:A(I)\to\ell^{\infty}_{2}(\mathbb{R})$ taking $f\in A(I)$ to (f(1),f(1))(f(-1),f(1)). Note that $A(I)$ has trivial involution, so all we need to check is that $\theta$ is multiplicative. However,
Because $\theta$ is bijective, this is a *-isomorphism. Although both $I$ and ${\mathcal{C}S}=K$ have the same complexification they are not nc affine homeomorphic, indeed their associated operator systems are $\ell^{\infty}_{2}(\mathbb{R})$ and $\mathbb{C}_{r}$, which are quite different. Indeed although $I={\rm Min}([-1,1])$ in the language of [8, Section 6], $K$ is not a Min nc convex set. Indeed $A(K)=\mathbb{C}_{r}$ is not an real OMIN operator system: it has nontrivial involution and is not a real function system in the technical sense of [9, 8].
We now explore $K$ and its complexification $K_{c}$ in detail, and use it to construct another example of a pure element $x\in K$ such that $x+i0$ is not pure in $K_{c}$.
Example 3.6.
Consider $K$ as defined above. From Theorem $3.1$ in [8], we know that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\mathrm{ncconv}_{\mathbb{R}}\{c(i)\}\}=\mathrm{ncconv}_{\mathbb{C}}\{c(i)\}$. A compression by $u$ shows $\pm i\in K_{c}$ so that $\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}\subseteq K_{c}$. Notice that we have
@@
so that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}$. Let $L=\mathrm{ncconv}_{\mathbb{C}}\{\pm 1\}$, the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. Because $\pm i\in i\cdot L$ we have that $K_{c}\subseteq iL$. Conversely, $\pm 1\in-i\cdot\mathrm{ncconv}\{\pm i\}$ so that $L\subseteq-iK_{c}$. This implies $iL\subseteq K_{c}$ and thus $iL=K_{c}$. However, $L=\mathrm{ncconv}\{\pm 1\}=\bigsqcup_{n}[-1,1]_{n}$ by Example $4.1$ in [43] or as can be deduced from Example 3.5. Also, $iL\cong L$ because multiplication by $i$ is a nc affine homeomorphism. Now, $L$ is compact, implying that $K_{c}$ is too, and thus $K$ is compact by Theorem 3.1. We have that $\pm 1\in L$ are the only nc extreme points also by Example $4.1$ in [43], so that $\pm c(i)$ are the only possible extreme points in $K$, and they are indeed extreme. Also, $0\in K_{c}$ is not pure as it is a convex combination of $\pm i$.
@@
For $K$ a real compact nc convex set, if $x\in K_{c}$ has $c(x)$ pure (resp. nc extreme, maximal) in $K$ then $x$ is pure (resp. nc extreme, maximal) in $K_{c}$.